# New versions of Frobenius and integral closure of ideals

### v2

@@ Abstract.
In this article, we define two new versions of integral and Frobenius closures of ideals which incorporate an auxiliary ideal and a real parameter. These additional ingredients are common in adjusting old definitions of ideal closures in order to generalize them to pairs, with an eye towards further applications in algebraic geometry. In the case of tight closure, similar generalizations exist due to N. Hara and K.I. Yoshida, as well as A. Vraciu. We study their basic properties and give computationally effective calculations of the adjusted tight, Frobenius, and integral closures in the case of affine semigroup rings in terms of the convex geometry of the associated exponent sets.
@@
Keyword: Integral closure, tight closure, Frobenius closure, semigroup rings, joint Hilbert-Kunz multiplicity
@@ 1. Introduction
In studying tight closure, one is naturally led to study test ideals, which uniformly witness tight closure relations for every ideal in the ring simultaneously. N. Hara [Har01] and K. E. Smith [Smi00] showed that test ideals in prime characteristic correspond in a natural way to multiplier ideals in birational geometry, and so are deeply related to geometric concepts like resolutions of singularities. The initial connection made by Hara and Smith was to the multiplier ideal of the trivial pair, but a strength of the multiplier ideal concept is that one may vary a $\mathbb{Q}$-divisor or an ideal of the ring and obtain information about the singularities in the locus of primes containing that ideal. For a survey of test ideals and their connections to multiplier ideals, we recommend [ST12].
Hara and K. Yoshida [HY03] introduced an adjusted form of tight closure, called $\mathfrak{a}^{t}$-tight closure (see 3.9 below), which incorporates an auxiliary ideal and a non-negative real parameter. The benefit to this adjustment is that now the appropriate notion of test ideals for $\mathfrak{a}^{t}$-tight closure corresponds to the multiplier ideal of the ideal pair $(R,\mathfrak{a}^{t})$. One disadvantage of the Hara-Yoshida tight closure is that it is not a true closure operation, in that the operation is not idempotent.
To rectify this issue, A. Vraciu [Vra08] introduced a new version of $\mathfrak{a}^{t}$-tight closure, which is a true ideal closure operation. Her version is contained in the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure, and captures subtly different information. In some cases the test ideals of the Hara-Yoshida and Vraciu versions agree, see [Vra08, Theorem 4.3]. Similar versions of adjusted tight and (sharp) Frobenius closures are studied in [STV17], where the authors use these closures to define and characterize $F$-rationality and sharp $F$-injectivity of the pair $(R,\mathfrak{a}^{t}),$ and also to define the ($F$-pure) threshold of the ideal $\mathfrak{a}.$
Our aim in this article is to repeat the adjustments of tight closure made by Hara-Yoshida and Vraciu to Frobenius and integral closure, which are two more classical closures near to the tight closure.
Definition (3.1, 3.12).
Fix a real number $t>0$, and let $I$ and $\mathfrak{a}$ be ideals in a ring $R$. Let $R^{\circ}$ all elements of $R$ not in a minimal prime of $R$. For the case of Frobenius closure, we assume that $R$ is of prime characteristic $p>0$.
@@
    $\overline{I}^{\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\right.\right\}.$
@@
    The Vraciu $\mathfrak{a}^{t}$-Integral closure of $I$, denoted ${}^{\mathfrak{a}^{t}}\overline{I}$, is defined by
    ${}^{\mathfrak{a}^{t}}\overline{I}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }n\gg 0,\,cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset\mathfrak{a}^{\lceil tn\rceil}I^{n}\right.\right\}.$
@@
    $I^{F\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
  4. (d)
@@
    ${}^{\mathfrak{a}^{t}}I^{F}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
Throughout Section 3, we show that the properties of these new adjusted closures are similar to the adjusted tight closures and their usual closures. In Subsection 3.4, we compare the new closures to each other and give an adjusted version of the Briançon-Skoda theorem.
Theorem (Theorem 3.19).
Let $\mathfrak{a},I$ be ideals in a ring $R$ of prime characteristic $p>0$ where $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are non-empty. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{\geq 0}$.
  1. (a)
  2. (b)
One unfortunate aspect of the adjusted closure operations are that they are even more difficult to compute than their usual counterparts. However, in some cases, they are related to colon ideals of the usual closure operation. Intriguingly, we also relate the adjusted tight and integral closure to the rational powers $\mathfrak{a}_{t}$ of an ideal (see 2.4).
Theorem (3.5, 3.10, Proposition 3.16).
Let $R$ be a ring and $I,\mathfrak{a}\subset R$ be ideals such that $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are non-empty. Further, let $t\geq 1$ be a rational number. Then,
@@
In Section 4, we carefully study the closures in the setting of affine semigroup rings. In particular, we are able to provide explicit calculations of the adjusted closures in terms of the convex geometry of the associated exponent sets of the ideals involved.
@@
Fix a real number $t>0$ and a field $k$ of prime characteristic $p>0$. Let $\sigma^{\vee}\subseteq\mathbb{R}^{d}$ be a rational polyhedral cone, $S=\sigma^{\vee}\cap\mathbb{Z}^{d}$ the semigroup of lattice points in $\sigma^{\vee}$, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$. In particular, there are finite subsets $\mathcal{A}$ and $\mathcal{B}$ of $S$ so that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ in $\mathbb{R}^{d}$. For a monomial $x^{\mathbf{v}}\in R$, we have
@@
These theorems show that in the affine semigroup ring case, the formulas which compute the $\mathfrak{a}^{t}$-tight and integral closures are independent of the characteristic of the underlying field. However, we show in an example (see Proposition 4.12) that the $\mathfrak{a}^{t}$-Frobenius closure can depend on the characteristic. We also show in the affine semigroup setting that the joint Hilbert-Kunz multiplicity introduced by Vraciu is computable via the volume of a certain polytope, which reflects the situation for usual Hilbert-Kunz multiplicity by work of K. Eto [Eto02, Theorem 2.2] and K-i. Watanabe [Wat00, Theorem 2.1].
@@
Fix a real number $t>0$ and a field $k$ of prime characteristic $p>0$. Let $S\subset\mathbb{N}^{d}$ be an affine semigroup, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$ of finite colength. Let $\mathcal{A}$ and $\mathcal{B}$ be finite subsets of $S$ such that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ and $\sigma^{\vee}$ for the cone defining $S$. If we set
@@
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{q\rightarrow\infty}\frac{1}{q^{d}}\ell\left(\frac{R}{\mathfrak{a}^{\lceil tq\rceil}I^{[q]}}\right)=\operatorname{vol}\overline{\mathcal{P}_{t}},$
@@
Finally, in order to demonstrate the theory to the fullest, we provide a thorough analysis of the adjusted closures in a specific affine semigroup ring, see Proposition 4.12. These examples demonstrate a “jumping-number” behavior for the adjusted closures in terms of the parameter $t$. For this example, we also compute the joint Hilbert-Kunz multiplicity as a function of the parameter $t$.
Acknowledgments: The authors are grateful to Rankeya Datta, Mark Johnson, Lance E. Miller, and Janet Vassilev for suggestions and comments. The first author was supported by grant number CEX2021-001142-S (Excelencia Severo Ochoa) funded by MICIU/AEI/10.13039/501100011033 (Ministerio de Ciencia, Innovación y Universidades, Spain).
@@ 2. Preliminaries
First, $0\in\mathbb{N}$ and $p$ will always denote a positive prime integer. All rings considered will be commutative with unit and Noetherian. For a ring $R$, we will let $R^{\circ}$ denote the multiplicative set of elements outside of the union of the minimal primes of $R$. That is, if
$S=\bigcup_{\mathfrak{p}\in\operatorname{Min}(R)}\mathfrak{p},$
then $R^{\circ}=R\setminus S$. For an ideal $I\subset R$, $\mu(I)$ denotes the minimal number of generators of $I$. Also, non-positive powers of ideals will be interpreted to mean the entire ring. Finally, for an $\mathbb{N}$-graded ring $R=\bigoplus_{n\geq 0}R_{n}$ and a real number $t$, we will use $R_{<t}$ (similarly $R_{\leq t}$, $R_{\geq t}$, $R_{>t}$) to denote the $R_{0}$-submodule of elements of $R$ of degree less than $t$.
We begin with a corollary to the usual Determinantal Trick for integral closure. While a similar result (without proof) appears in [Vas05, Proposition 1.58] for integrally closed ideals, we provide one for completeness. The ideal $IJ:_{R}J$ described below is called the $J$-basically full closure of $I$. See [ERGV23, Section 4] for a modern treatment.
Theorem 2.1 (The Determinantal Trick).
If $I$ and $J$ are ideals in a ring $R$ with $J\cap R^{\circ}\neq\varnothing$, then $(IJ:_{R}J)\subset\overline{I}$, the integral closure of $I$.
@@
Thus, there exists $m\in\mathbb{N}$ such that
$0=(x^{n}+r_{1}x^{n-1}+\cdots+r_{n})^{m}=x^{mn}+s_{1}x^{mn-1}+\cdots+s_{mn}$
where each $s_{j}$ is a sum of terms of the form $r_{i_{1}}r_{i_{2}}\cdots r_{i_{m}}$ with $i_{1}+i_{2}+\cdots+i_{m}=j$. Hence $s_{j}\in I^{j}$. This shows that $x\in\overline{I}$. ∎
@@
    $I^{F}=\left\{x\in R\left|\text{ for some }e\in\mathbb{N},\,x^{p^{e}}\in I^{\left[p^{e}\right]}\right.\right\}.$
@@
    $I^{*}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,\,cx^{p^{e}}\in I^{\left[p^{e}\right]}\right.\right\}.$
For an ideal $I$ in a ring $R$ of prime characteristic $p$, we have that $I\subset I^{F}\subset I^{*}\subset\overline{I}$, but each of these containments are typically strict. However, in a regular ring, $I=I^{*}$ for all $I$ by [HH90, Theorem 4.4], so many singularity types in prime characteristic are defined in terms of whether some or all ideals are closed with respect to these operations. For example, a ring is called weakly $F$-regular if $I^{*}=I$ for all ideals $I$.
The computation of tight and integral closure reduce readily to domains (see [HH90, Proposition 6.25.(a)] and [SH06, Proposition 1.1.5]) as elements of the ring $R$ are in the tight or integral closure of an ideal $I$ if and only if they are in the tight or integral closure of $I+\mathfrak{p}/\mathfrak{p}$ in $R/\mathfrak{p}$ where $\mathfrak{p}$ ranges over all minimal primes of $R$.
It is well-known that the Frobenius closure does not exhibit this property, that is, the inclusion
$I^{F}\subset\left\{x\in R\mid x+\mathfrak{p}\in(I+\mathfrak{p})^{F}/\mathfrak{p}\text{ for any }\mathfrak{p}\in\operatorname{Min}R\right\}$
may be proper; we provide an example below.
@@
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$. One may show that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and that $z\in(x)^{*}\setminus(x)^{F}.$ Indeed, if $z\in(x)^{F},$ then $z^{3^{e}}\in I^{[3^{e}]}=(x^{3^{e}})$ for some $e.$ Write $13^{e}=2k+$ 1 for some $k.$ Then
@@
Let $R$ be a ring, $\mathfrak{a}\subset R$ be an ideal, and $t=r/s\in\mathbb{Q}_{\geq 0}$. Then, the $t$-th rational power of $\mathfrak{a}$, denoted $\mathfrak{a}_{t}$, is defined by $\mathfrak{a}_{t}=\left\{x\in R\left|\,x^{s}\in\overline{\mathfrak{a}^{r}}\right.\right\}.$
Despite being introduced many years ago, the rational powers of ideals have received little attention until recently. See [SH06, Section 10.5] for a treatment of rational powers. For some modern results utilizing them, see [Ciu20] and [BDHM24].

### v3

@@ Abstract.
In this article, we define three new operations on ideals which generalize integral closure and Frobenius closure of ideals, whose definitions incorporate an auxiliary ideal and a real parameter. These additional ingredients are common in adjusting old definitions of ideal closures in order to generalize them to pairs, with an eye towards further applications in algebraic geometry. In the case of tight closure, similar generalizations exist due to N. Hara and K.I. Yoshida, as well as A. Vraciu, and in the case of Frobenius closure, to K. Schwede. We study their basic properties and give computationally effective calculations of the adjusted tight, Frobenius, and integral closures in the case of affine semigroup rings in terms of the convex geometry of the associated exponent sets. Finally, as applications, we study submodules of the fraction field of a domain defined in terms of our adjusted closures and an $F$-nilpotent property for pairs.
@@
Keyword: Integral closure, tight closure, Frobenius closure, semigroup rings, joint Hilbert-Kunz multiplicity, F-nilpotence of pairs
@@ 1. Introduction
In studying tight closure, one is naturally led to study test ideals, which uniformly witness tight closure relations for every ideal in the ring simultaneously. N. Hara [Har01] and K. E. Smith [Smi00] showed that test ideals in prime characteristic correspond in a natural way to multiplier ideals in birational geometry, and so are deeply connected to geometric concepts like resolutions of singularities. The initial correspondence made by Hara and Smith was to the multiplier ideal of the trivial pair, but a strength of the theory of multiplier ideals is that one may vary an ideal of the ring and obtain information about the singularities in the locus of primes containing that ideal. For a survey of test ideals and their connections to multiplier ideals, we recommend [ST12].
Hara and K.I. Yoshida [HY03] introduced an adjusted form of tight closure, called $\mathfrak{a}^{t}$-tight closure (see 3.11 below), which incorporates an auxiliary ideal and a real parameter. The benefit to this adjustment is that now the appropriate notion of test ideals for $\mathfrak{a}^{t}$-tight closure corresponds to the multiplier ideal of the ideal pair $(R,\mathfrak{a}^{t})$. One disadvantage of the Hara-Yoshida tight closure is that it is not an ideal closure operation, in that the operation is not idempotent. [^1] We would here like to caution the reader that, throughout the article, we follow the slightly unfortunate convention of referring to the Hara-Yoshida style operations as “closures” even though they are only pre-closures, see Remark 2.2.
To rectify this issue, A. Vraciu [Vra08] introduced a new version of $\mathfrak{a}^{t}$-tight closure, which is a true ideal closure operation. Her version is contained in the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure, and captures subtly different information. In some cases the test ideals of the Hara-Yoshida and Vraciu versions agree, see [Vra08, Theorem 4.3]. Vraciu also introduced a Hilbert-Kunz-like multiplicity which detects equality of her $\mathfrak{a}^{t}$-tight closure among other properties. Our aim in this article is to repeat the adjustments of tight closure made by Hara-Yoshida and Vraciu to Frobenius and integral closure, which are two more classical closures near to the tight closure.
Definition (3.1, 3.14).
Fix a real number $t\in\mathbb{R}_{>0}$, and let $I$ and $\mathfrak{a}$ be ideals in a ring $R$, which we typically assume are of positive height. Let $R^{\circ}$ be the set of all elements of $R$ not in any minimal prime of $R$. For the case of Frobenius closure, we assume that $R$ is of prime characteristic $p>0$.
@@
[…]
@@
[…]
@@
[…]
Though it seems natural to include a Vraciu version of integral closure alongside the Hara-Yoshida version, we immediately show that the obvious definition coincides with the usual integral closure (see 3.3). The diagram below illustrates the relationships among these operations.
[…]
In [Sch08], K. Schwede provided a similar adjustment to the Frobenius closure to study $F$-purity of pairs $(R,\mathfrak{a}^{t})$. Our version is larger than his in general, due to a slightly larger exponent on the auxiliary ideal, and we provide examples to compare them in Section 4.2. In contrast, we also show our definitions for adjusted tight and integral closure are insensitive to certain linear changes in the exponent on the auxiliary ideal (see Theorem 4.9).
Throughout this article, we show that the properties of these new adjusted closures are similar to both their unadjusted versions as well as the adjusted tight closures of Hara-Yoshida and Vraciu. One unfortunate aspect of the adjusted closure operations are that they are even more difficult to compute than their usual counterparts. However, in some cases, they are related to colon ideals of the usual closure operation. Intriguingly, we also relate the adjusted tight and integral closure to the rational powers $\mathfrak{a}_{t}$ of an ideal (see 2.7).
Theorem (3.5, 3.12, Proposition 3.18).
Let $R$ be a ring and $I,\mathfrak{a}\subset R$ be ideals of positive height. Further, let $t>1$ be a rational number. Then,
@@
We also demonstrate a parallel between the valuation theory involved in the computation of ordinary integral closure and the Hara-Yoshida integral closure which we define.
Theorem (Theorem 3.9).
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ ideals of positive height, and $t\in\mathbb{R}_{>0}$. Let $\mathcal{RV}(I)$ be the set of Rees valuations of $I$. An element $x\in R$ is in $\overline{I}^{\mathfrak{a}^{t}}$ if and only if for all $\nu\in\mathcal{RV}(I)$, $\nu(x)\geq\nu(I)-t\nu(\mathfrak{a})$.
In Section 4, we compare the new closure operations to each other in detail. Perhaps of most interest in this section, we develop an $\mathfrak{a}^{t}$-analogue for the tight closure Briançon-Skoda theorem of Hochster-Huneke ([HH90, Thm. 5.4]), which controls the integral closure of powers of an ideal by fitting them inside the tight closure of a smaller power of the same ideal. In [Hun92], Huneke also provided a uniform Briançon-Skoda theorem which, under some mild hypotheses on the ring, makes uniform the shifted exponent. We obtain a similar uniform tight closure Briançon-Skoda theorem for our adjusted closure operations.
Theorem (Theorem 4.5).
Let $\mathfrak{a},I$ be ideals of positive height in a ring $R$ of prime characteristic $p>0$. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{>0}$.
  1. (a)
  2. (b)
Further, suppose $(R,\mathfrak{m})$ is local of dimension $\dim R=n$ and $R/\mathfrak{m}$ is infinite.
  1. (c)
    Then $\overline{I^{n+m}}^{\mathfrak{a}^{t}}\subset(I^{m+1})^{*\mathfrak{a}^{t}}$ for all $m\geq 0$, and
  2. (d)
    $\overline{I^{n+m}}^{\mathfrak{a}^{t}}\subset(I^{m})^{F\mathfrak{a}^{t}}$ for all $m\geq 1$.
We also provide a graded version of the Briançon-Skoda theorem for the adjusted tight closure (see Proposition 4.14) along the lines of [Smi97], which analyzes which graded components of the ring which are forced into the tight closure of an ideal primary to the homogeneous maximal ideal.
As an application of our results, we generalize another collection of classical definitions to the pairs setting with our new closures. In particular, it is elementary to show that the normalization $\overline{R}$ of a domain $R$ is the set of fractions $x/y$ in the fraction field of $R$ such that $x\in\overline{(y)}$. Similarly, the weak normalization $^{*}R$ of a domain $R$ is the set of fractions $x/y$ such that $x\in(y)^{F}$. With these perspectives in hand, we define the following.
Definition.
Let $R$ be a domain, $t\in\mathbb{R}_{>0}$, and let $\mathfrak{a}$ be an ideal of positive height in $R$. Then, we define the Hara-Yoshida $\mathfrak{a}^{t}$-normalization of $R$ to be
[…]
Analogously, we also define the Hara-Yoshida $\mathfrak{a}^{t}$-weak normalization $R^{F\mathfrak{a}^{t}}$, and the Vraciu $\mathfrak{a}^{t}$-weak normalization $^{\mathfrak{a}^{t}}R^{F}$ for domains of prime characteristic $p>0$.
Because the adjusted tight closures of principal ideals agree with the corresponding integral closures, we do not have need to define similar objects involving the tight closures. We demonstrate that membership in these sets is well-defined, and further, that each forms an $R$-submodule of the fraction field (5.2), and indeed, $^{\mathfrak{a}^{t}}R^{F}$ is even a subring (5.4) of $\overline{R}.$ Through 5.6 and 5.7, we show that each of the natural inclusions in the diagram below can be strict.
$R{{}^{*}R}$$^{\mathfrak{a}^{t}}R^{F}$$R^{F\mathfrak{a}^{t}}$$\overline{R}$$\overline{R}^{\mathfrak{a}^{t}}$}
Ideal closures in prime characteristic are intimately related to $F$-singularities, which are singularities defined in terms of the Frobenius map on a ring of prime characteristic. For details about $F$-singularities in general, we direct the reader to recent surveys by Schwede-Smith [SS25] and L. Ma-T. Polstra [MP25]. One recently-defined $F$-singularity of interest is $F$-nilpotence (see 5.10), defined by V. Srinivas-S. Takagi in [ST17], which is defined in terms of the Frobenius action on local cohomology, or equivalently (under some mild hypotheses) to equality of tight and Frobenius closure of parameter ideals (see [PQ19]).
Just as the Hara-Yoshida tight closure and Schwede’s sharp $F$-closure have been used to study $F$-singularities of pairs, we use the Hara-Yoshida tight closure to define $F$-nilpotence of pairs.
Definition.
Let $(R,\mathfrak{m})$ be an excellent, equidimensional, reduced local ring, $t\in\mathbb{R}_{>0}$, and suppose $\mathfrak{a}$ is an ideal of $R$ of positive height. We say the pair $(R,\mathfrak{a}^{t})$ is $F$-nilpotent if $\mathfrak{q}^{*\mathfrak{a}^{t}}=\mathfrak{q}^{F}$ for all ideals $\mathfrak{q}\subset R$ generated by a system of parameters for $R$.
Alongside studying basic properties of this new singularity of pairs, we also demonstrate an equivalence of the definition given above to a condition involving the top local cohomology module in the case that the ring is Cohen-Macaulay, although we expect the theorem to hold more generally.
Theorem (Theorem 5.16).
Let $(R,\mathfrak{m})$ be an excellent, Cohen-Macaulay local domain of prime characteristic $p>0$. Then, $(R,\mathfrak{a}^{t})$ is $F$-nilpotent if and only if $0^{*\mathfrak{a}^{t}}_{H^{d}_{\mathfrak{m}}(R)}=0^{F}_{H^{d}_{\mathfrak{m}}(R)}$.
@@
In Section 6, we carefully study the adjusted closure operations in the setting of monomial ideals in affine semigroup rings. In particular, we are able to provide explicit calculations of the adjusted closures in terms of the convex geometry of the associated exponent sets of the ideals involved, which has the added effect of providing concrete examples of the Hara-Yoshida and Vraciu tight closure which were otherwise missing from the literature. These results also extend well-known results for computing the ordinary integral closure and Hilbert-Kunz multiplicity.
@@
Fix $t\in\mathbb{R}_{>0}$ and a field $k$ of prime characteristic $p>0$. Let $\sigma^{\vee}\subset\mathbb{R}^{d}$ be a rational polyhedral cone, $S=\sigma^{\vee}\cap\mathbb{Z}^{d}$ the semigroup of lattice points in $\sigma^{\vee}$, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$. In particular, there are finite subsets $\mathcal{A}$ and $\mathcal{B}$ of $S$ so that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ in $\mathbb{R}^{d}$. For a monomial $x^{\mathbf{v}}\in R$, we have
@@
These theorems show that in the affine semigroup ring case, the formulas which compute the $\mathfrak{a}^{t}$-tight and integral closures are independent of the characteristic of the underlying field. However, we show in an example (see 6.10) that the Vraciu $\mathfrak{a}^{t}$-Frobenius closure can depend on the characteristic. We also show in the affine semigroup setting that the joint Hilbert-Kunz multiplicity introduced by Vraciu is computable via the volume of a certain polytope, which reflects the situation for usual Hilbert-Kunz multiplicity by work of K. Eto [Eto02, Theorem 2.2] and K-i. Watanabe [Wat00, Theorem 2.1].
@@
Fix $t\in\mathbb{R}_{>0}$ and a field $k$ of prime characteristic $p>0$. Let $S\subset\mathbb{N}^{d}$ be an affine semigroup, $R=k[S]$, and let $I$ and $\mathfrak{a}$ be monomial ideals in $R$ of finite colength. Let $\mathcal{A}$ and $\mathcal{B}$ be finite subsets of $S$ such that $I=(x^{\mathbf{b}}\mid\mathbf{b}\in\mathcal{B})$ and $\mathfrak{a}=(x^{\mathbf{a}}\mid\mathbf{a}\in\mathcal{A})$. We write $H$ for the convex hull of $\mathcal{A}$ and $\sigma^{\vee}$ for the cone defining $S$. If we set
@@
$e_{HK}(\mathfrak{a}^{t};I)=\lim\limits_{q\rightarrow\infty}\frac{1}{p^{ed}}\lambda_{R}\left(\frac{R}{\mathfrak{a}^{\lceil tq\rceil}I^{[p^{e}]}}\right)=\operatorname{vol}\overline{\mathcal{P}_{t}},$
@@
Finally, in order to demonstrate the theory to the fullest, we provide a thorough analysis of the adjusted closures in a specific affine semigroup ring, see 6.10. These examples demonstrate discrete shifts in the adjusted closures in finitely many intervals for the parameter $t$. For this example, we also compute the joint Hilbert-Kunz multiplicity as a function of $t$.
Acknowledgements.
The authors are grateful to Rankeya Datta, Mark Johnson, Lance E. Miller, Janet Vassilev, and Thomas Polstra for useful conversations. Furthermore, we want to thank Kevin Tucker and Yairon Cid-Ruiz for suggestions about the uniform Briançon-Skoda theorem. The first author was supported by grant number CEX2021-001142-S (Excelencia Severo Ochoa) funded by MICIU/AEI/10.13039/501100011033 (Ministerio de Ciencia, Innovación y Universidades, Spain).
@@ 2. Preliminaries
First, $0\in\mathbb{N}$ and $p$ will always denote a positive prime integer. All rings are assumed to be commutative with unit and noetherian. For a ring $R$, we will let $R^{\circ}$ denote the multiplicative set of elements outside of the union of the minimal primes of $R$. That is, $R^{\circ}=R\setminus\bigcup_{\mathfrak{p}\in\operatorname{Min}(R)}\mathfrak{p}$. For an ideal $I\subset R$, $\mu(I)$ denotes the minimal number of generators of $I$. We will use $\lambda_{R}(\bullet)$ for the length of an $R$-module. Also, non-positive powers of ideals will be interpreted as the entire ring.
For an $\mathbb{N}$-graded ring $R=\bigoplus_{n\geq 0}R_{n}$ and a real number $t$, we will use $R_{<t}$ to denote the $R_{0}$-submodule generated by elements of $R$ of degree less than $t$ (similarly $R_{\leq t}$, $R_{\geq t}$, $R_{>t}$). We will slightly overload this notation for the usual number rings $\mathbb{Z}$, $\mathbb{Q}$, $\mathbb{R}$, e.g. using $\mathbb{R}_{>0}$ for the set of positive real numbers.
Our article primarily concerns ideal closure operations.
Definition 2.1.
Let $R$ be a ring and let $I\subset R$ be an ideal. We will call a function from the set of ideals of $R$ to itself an operation on ideals. An operation on ideals $\bullet^{c}$ is called a pre-closure if
  1. (a)
    $\bullet^{c}$ is extensive, that is, $I\subset I^{c}$, and
  2. (b)
    $\bullet^{c}$ is order-preserving, that is, $I\subset J$ implies $I^{c}\subset J^{c}$.
If a preclosure $\bullet^{c}$ additionally satisfies
  1. (c)
    $\bullet^{c}$ is idempotent, that is, $(I^{c})^{c}=I^{c}$,
then we say $\bullet^{c}$ is an ideal closure operation.
We will soon survey the definitions of several important ideal closure operations in rings of prime characteristic $p>0$.
Remark 2.2.
Even though several operations on ideals we consider later are simply pre-closures, due to prior authors’ convention, they are commonly referred to as closures. We will indicate when defining an operation whether it is a pre-closure or closure, but will typically refer to both kinds of operations as “closure operations” in the text.
We begin with a corollary to the usual Determinantal Trick for integral closure. While a similar result (without proof) appears in [Vas05, Proposition 1.58] for integrally closed ideals, we provide one for completeness. The ideal $IJ:_{R}J$ studied below is called the $J$-basically full closure of $I$, so one can summarize the following theorem by saying the $J$-basically full closure of $I$ is contained in its integral closure. See [ERGV23, Section 4] for a modern treatment of the basically fully closure.
Theorem 2.3.
If $R$ is a ring and $I,J\subset R$ are ideals of positive height, then $(IJ:_{R}J)\subset\overline{I}$, the integral closure of $I$.
@@
Thus, there exists $m\in\mathbb{N}$ such that $0=(x^{n}+r_{1}x^{n-1}+\cdots+r_{n})^{m}=x^{mn}+s_{1}x^{mn-1}+\cdots+s_{mn}$ where each $s_{j}$ is a sum of terms of the form $r_{i_{1}}r_{i_{2}}\cdots r_{i_{m}}$ with $i_{1}+i_{2}+\cdots+i_{m}=j$. Hence $s_{j}\in I^{j}$. This shows that $x\in\overline{I}$. ∎
@@
[…]
@@
[…]
Both Frobenius and tight closure are ideal closure operations.
For an ideal $I$ in a ring $R$ of prime characteristic $p>0$, we have that $I\subset I^{F}\subset I^{*}\subset\overline{I}$, but each of these containments are typically strict. However, in a regular ring, $I=I^{*}$ for all $I$ by [HH90, Theorem 4.4], so many singularity types in prime characteristic are defined in terms of whether some or all ideals are closed with respect to these operations. For example, a ring is called weakly $F$-regular if $I^{*}=I$ for all ideals $I$.
We will later discuss $F$-nilpotent rings, which are another singularity type that can be defined in terms of ideal closures. Initially studied by Blickle-Bondu in [BB05] under the name close to $F$-rational, $F$-nilpotent rings were systematically studied by Srinivas-Takagi in [ST17], who showed that low-dimensional rings in characteristic zero which are of $F$-nilpotent type satisfy certain Hodge-theoretic conditions related to the Weak Ordinarity conjecture. The initial definition of $F$-nilpotent rings involves the canonical Frobenius action on local cohomology, the technicalities of which we avoid here, but Polstra-Quy showed in [PQ19] that the following definition is equivalent to the one given by Blicke-Bondu and Srinivas-Takagi in the case that $R$ is excellent and equidimensional.
Definition 2.5.
Let $(R,\mathfrak{m})$ be an excellent, equidimensional local ring. Then, we say that $R$ is $F$-nilpotent if $\mathfrak{q}^{*}=\mathfrak{q}^{F}$ for all ideals $\mathfrak{q}$ of $R$ generated by a system of parameters.
The computation of tight and integral closure reduce readily to domains (see [HH90, Proposition 6.25.(a)] and [SH06, Proposition 1.1.5]) as elements of the ring $R$ are in the tight or integral closure of an ideal $I$ if and only if they are in the tight or integral closure of $(I+\mathfrak{p})/\mathfrak{p}$ in $R/\mathfrak{p}$ where $\mathfrak{p}$ ranges over all minimal primes of $R$ ($\operatorname{Min}R$). It is well-known that the Frobenius closure does not exhibit this property, that is, the inclusion
$I^{F}\subset\left\{x\in R\mid x+\mathfrak{p}\in(I+\mathfrak{p})^{F}/\mathfrak{p}\text{ for all }\mathfrak{p}\in\operatorname{Min}R\right\}$
may be proper; we provide an explicit example below.
@@
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$. It is easy to check that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and that $z\in(x)^{*}\setminus(x)^{F}.$ Indeed, if $z\in(x)^{F},$ then $z^{3^{e}}\in I^{[3^{e}]}=(x^{3^{e}})$ for some $e.$ Write $3^{e}=2k+1$ for some $k.$ Then
@@
[…]
Despite being introduced over fifty years ago, the rational powers of ideals have unfortunately received little attention until recently. See [SH06, Section 10.5] for a treatment of rational powers. For some modern results utilizing them, see [Ciu20] and [BDHM24].